# Systematic Generalization in Language Models Scales with Information Entropy

### v1

@@ 4.6 Results and Discussion
We argue that these results indicate that the Transformer is more information efficient than the RNN and CNN under full support, but that the models are similar under lacking support. In practice, this means that the Transformer is able to generalize to novel uses of component parts, as long as the overall syntactic pattern—constituted by the conjunctive phrase in our experiments—is frequent in the training data. This interpretation aligns with Loula et al. (2018), showcasing how our framework can provide a theoretical explanation for previous results.
We note that even though all models can generalize under high entropy for both experiments, their performance is subpar for low entropy, e.g. $\mathrm{H=1}$. This indicates a rather high lower bound on the entropy required for systematic generalization, both for vertical and horizontal scaling. This shows that systematic generalization remains an open problem when there is no strict architectural prior, and it remains unclear whether or not it is possible to build a model that performs well on low entropy without such priors.
4.6.1 Supplemental results
@@
In Figure 6(b), we report the performance of different positional encoding types. Although all encoding types achieve near-ceiling performance at high entropy, absolute encoding performs better for all levels of $\mathrm{H}$. From our task formulation in Section 3, it is clear that positional information is crucial for solving the tasks. We attribute the efficiency of absolute encodings to the fact that the positional information of the verbs is independent of the remainder of the sequence. This invariance has to be learned by the other embedding types, making the absolute variant more information efficient.
@@ 5 Previous Work
Most similar to our work is Zhou et al. (2023), who also studies how aspects of the SCAN grammar affects compositional generalization. We see compositionality as a set of primitives composed by a composition function, and our work focuses on how models capture this composition function depending on properties of the primitives. Their work, however, do not make this distinction, as their complexification process modify both the composition function and the distribution of primitives. The main distinction is hence that our work relates compositionality to dataset-level regularities, while they relate it to sample-level features.
In another domain, Wiedemer et al. (2023) approaches systematicity in an image generation problem by factorizing the latent variables of the input image. The authors show that the support of these variables must be the same in the training and test distributions for any model to generalize systematically. We take a more fine-grained approach, and our results demonstrate that support alone is insufficient as a predictor of systematic generalization; the performance also depends on the entropy, and higher levels of entropy in the training distribution can compensate for lower support.
There is little existing research on how to measure systematicity from a data-centric perspective. However, there is ongoing work on formalizing compositionality in terms of data complexity (Elmoznino et al., 2024) and kernel-theory (Lippl and Stachenfeld, 2024). Most similar to our approach is Keysers et al. (2020), which also quantifies the level of systematic generalization using distributional information. Specifically, they quantify the compositionality of an experiment as the divergence between the distribution of component parts and combinations in the training and test data. Their work differs from ours since they focus on the case where the distribution of component parts is as similar as possible between the training and test data, while our work focuses on the case where the distribution of component parts is inverted from the training to the test set for some parts of the input.
@@ 6 Conclusion
In this work, we show that the performance of sequence-to-sequence models scales with entropy. We also show that this scaling is independent of the number of unique samples. Contrary to previous work, our findings demonstrate that these models are capable of systematic generalization. However, this requires high entropy. This poses a new question: Is it possible to achieve systematicity also at low entropy without encoding explicit architectural priors? Through our formalizations, we hope to facilitate future work that tries to answer this question, providing a clear method for assessing progress in systematic generalization research.
@@
In this work we have considered an aspect of systematicity that is concerned with embedded sentences as defined by Hadley (1994). There is no guarantee that our findings generalize to scenarios that are different from ours, but that still falls in under some definition of systematic generalization.
@@ Appendix C Computational budget
  • Permutation equivariant sweep: 24 hours

### v2

@@ 4.6 Results and Discussion
We argue that these results indicate that the Transformer is more information-efficient than the RNN and CNN under full support, but that the models are similar under lacking support. In practice, this means that the Transformer is quicker to generalize to novel uses of component parts, as long as the overall syntactic pattern—constituted by the conjunctive phrase in our experiments—is frequent in the training data. This interpretation aligns with the findings of Loula et al. (2018), showcasing how our framework can provide a theoretical explanation for previous results.
We note that even though all models can generalize under high entropy for both experiments, their performance is subpar for low entropy, e.g. $\mathrm{H}=1$. This indicates a rather high lower bound on the entropy required for systematic generalization, both for vertical and horizontal scaling, and it shows that all models are unable to meet requirement iii from Definition 3. Furthermore, our results show that systematic generalization remains an open problem when there is no prior, and it remains unclear whether or not it is possible to build a model that performs well on low entropy without architectural priors as in Gordon et al. (2019) or priors in the training procedure as in Lake and Baroni (2023).
4.6.1 Supplemental results
@@
In Figure 6(b), we report the performance of different positional encoding types. Although all encoding types achieve near-ceiling performance at high entropy, absolute encoding performs better for all levels of $\mathrm{H}$. From our task formulation in Section 3, it is clear that positional information is crucial for solving the tasks. We attribute the efficiency of absolute encodings to the fact that the positional information of the verbs is independent of the remainder of the sequence. This invariance has to be learned by the other embedding types, making the absolute variant more information-efficient.
@@ 5 Previous Work
The work most similar to ours is Zhou et al. (2023), who also studies how data-centric aspects affect compositional generalization. In our work, we see compositionality as characterized by a set of primitives composed by a composition function, and we focus on how models capture this composition function depending on properties of the primitives. Their work, however, does not make this distinction, as their complexification process modify both the composition function and the distribution of primitives. The main distinction is hence that our work relates compositionality to dataset-level regularities, while they relate it to sample-level features.
Our work is also similar to Keysers et al. (2020), who quantifies the level of systematic generalization using distributional information. Specifically, they quantify the compositionality of an experiment as the divergence between the distribution of component parts and combinations in the training and test data. Their work differs from ours since they focus on the case where the distribution of component parts is as similar as possible between the training and test data, while our work focuses on the case where the distribution of component parts is inverted from the training to the test set for some parts of the input.
In another domain, Wiedemer et al. (2023) approaches systematicity in an image generation problem by factorizing the latent variables of the input image. The authors show that the support of these variables must be the same in the training and test distributions for any model to generalize systematically. We take a more fine-grained approach and our results demonstrate that support alone is insufficient as a predictor of systematic generalization; the performance also depends on the entropy, and higher levels of entropy in the training distribution can compensate for lower support.
There is also recent work on formalizing compositionality in terms of data complexity (Elmoznino et al., 2024), kernel-theory (Lippl and Stachenfeld, 2025), using graph formalisms (Ram et al., 2024), and decomposition strategies (Jarvis et al., 2023). Our work is similar to these in the sense that we try to formalize some properties of systematic generalization, which is closely tied to compositionality.
@@ 6 Conclusion
In this work, we show that systematic generalization of sequence-to-sequence models scales with information entropy. Through supporting experiments, we also show that this relationship is independent of the number of unique samples seen during training. Contrary to previous work, our findings demonstrate that these models are capable of systematic generalization even without any built-in architectural priors that incentivize compositional solutions. However, this requires that the training data have high entropy w.r.t. the parts that the target composition function operates on. This poses a new question: Is it possible to achieve systematicity also at low entropy without compositional priors? Through our formalizations, we hope to facilitate future work that attempts to answer this question, providing a clear method for assessing progress in systematic generalization research.
@@
In this work, we have considered an aspect of systematicity that is concerned with embedded sentences as defined by Hadley (1994). There is no guarantee that our findings generalize to scenarios that are different from ours, but that still falls in under some definition of systematic generalization.
@@
Acknowledgements
We acknowledge Sigma2—the National Infrastructure for High-Performance Computing and Data Storage in Norway—for providing access to the LUMI super-computer, part of the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium. We also want to thank Lilja Øvrelid, Erik Velldal, and David Samuel at the Language Technology Group for discussions and feedback on the manuscript.
@@ Appendix C Computational budget
  • Permutation equivariant sweep: 24 hours.